If you are going 80 mph (speed) and have to travel 503 miles what is your arrival time?
Found 2 solutions by jim_thompson5910, Alan3354:
Answer by jim_thompson5910(35256)

d = rt

503 = 80t

503/80 = t

6.2875 = t

t = 6.2875


So the arrival time is 6.2875 hours

Answer by Alan3354(69443)

If you are going 80 mph (speed) and have to travel 503 miles what is your arrival time?
-------------
d = r*t
t = d/r = 503/80 hours enroute
The arrival time depends on the start time.